OpenStudy (anonymous):

2 log base 3 sq.root. 27 ?!?!!?!?!?!!??!

OpenStudy (anonymous):

2 x logbase3(sqrt(27))?

OpenStudy (anonymous):

\[2 \log _{3} \sqrt{27} \]

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

log base 3 of radical 27 is basicaly saying 3 to the what power = sqrt(27)

OpenStudy (anonymous):

and i don't understand logs in general

OpenStudy (anonymous):

but theres a 2 in front..

OpenStudy (anonymous):

you multiply by two after you get the first part

OpenStudy (anonymous):

its just multiplication in front

OpenStudy (anonymous):

do u mind drawing this out ? sorry I'm really stupid :(

OpenStudy (ash2326):

\[2 log_{3} (\sqrt {27})\] x=log y to the base b means b^x=y

OpenStudy (anonymous):

we actually just did this stuff today and yesterday in math class

OpenStudy (anonymous):

uh oh...maybe you're in my math class lol ...

OpenStudy (ash2326):

and we have the property c log y= log y(^c)

OpenStudy (anonymous):

lol, do you have algebra 2 im guessing?

OpenStudy (anonymous):

nope...im still very confused about the 2 in front though...this whole thing makes no sense to me

OpenStudy (anonymous):

the two in front is just multiplication, its like y = 2x, the two is multiplied

OpenStudy (ash2326):

now \( 2 \log_{3} (\sqrt{27})\) \(\log_{3}{(\sqrt{27})}^2\) \(\log_{3} 27\) we know that 3^3=27 so it's \(\log_{3} 27=3\)

OpenStudy (anonymous):

not sure what the guy above me is doing. How are you squaring the radical 27?

OpenStudy (anonymous):

i am literally so confused

OpenStudy (anonymous):

well, whatever he is doing is right.

OpenStudy (anonymous):

the answer is 3

OpenStudy (ash2326):